OpenStudy (maxdrizner):

Barry has been watching the geese that live in his neighborhood. The number of geese changes each week. n f(n) 1 56 2 28 3 14 4 7 Which function best shows the relationship between n and f(n)? (1 point)

OpenStudy (maxdrizner):

f(n) = 28(0.5)n f(n) = 56(0.5)n-1 f(n) = 56(0.5)n f(n) = 112(0.5)n-1

OpenStudy (anonymous):

right

OpenStudy (maxdrizner):

the answers are in the first comment

OpenStudy (anonymous):

i dont know this i was just sent her

OpenStudy (maxdrizner):

ok

OpenStudy (anonymous):

Its c

OpenStudy (maxdrizner):

i thought you didnt know how to do this

OpenStudy (anonymous):

That was a joke

OpenStudy (maxdrizner):

mhhmm

OpenStudy (anonymous):

But really its c

OpenStudy (anonymous):

i know this one my Butt told me

OpenStudy (maxdrizner):

i hope its right

OpenStudy (anonymous):

it came from end to another

OpenStudy (maxdrizner):

so now you are talking shet

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

what does a medal do for you

OpenStudy (fatimab2000):

its B